# Cortex AISQL: A Production SQL Engine for Unstructured Data

### v2

@@ Abstract.
Integrating AI’s semantic capabilities directly into SQL is a key goal, as it would enable users to write declarative queries that seamlessly blend relational operations with semantic reasoning. This would allow for effortless querying of both structured and unstructured data. However, achieving production-scale efficiency for semantic operations presents significant hurdles. Semantic operations inherently cost more than traditional SQL operations and possess different latency and throughput characteristics, making their application at large scales prohibitively expensive. Moreover, current query execution engines are not designed to optimize semantic operations, often leading to suboptimal query plans. This optimization challenge is compounded because the cost and selectivity of semantic operations are typically unknown at the time of compilation. Snowflake’s Cortex AISQL is a query execution engine that addresses these challenges through three novel techniques informed by production deployment data from Snowflake customers. First, AI-aware query optimization treats AI inference cost as a first-class optimization objective, reasoning about large language model (LLM) cost directly during query planning to achieve 2–8 speedups. Second, adaptive model cascades reduce inference costs by routing most rows through a fast proxy model while escalating uncertain cases to a powerful oracle model, achieving 2–6 speedups while maintaining 90–95% of oracle model quality. Third, semantic join query rewriting lowers the quadratic time complexity of join operations to linear through reformulation as multi-label classification tasks, achieving 15–70 speedups with often improved prediction quality. AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding.
@@ 1. Introduction
The growth of large language models (LLMs) (brown2020language; ouyang2022training; touvron2023llama; wei2022chain) has transformed how organizations interact with data. Modern enterprises now store vast volumes of unstructured content – documents, images, audio, and text – alongside traditional structured data in relational tables. While contemporary data warehouses excel at processing structured data with SQL, they lack native support for semantic reasoning over unstructured information. As a result, users who need to filter documents by topic relevance, join tables based on semantic similarity, or summarize customer feedback must export data to external systems, write custom scripts, or build complex pipelines that orchestrate LLM APIs outside the database. The result is significant friction: data must be moved across systems, increasing latency and cost; manual orchestration becomes error-prone; and the database loses the ability to optimize the workload end-to-end.
@@
We present Snowflake’s Cortex AISQL, a SQL engine that enables users to talk to their data, both structured and unstructured, through native AI operators integrated directly into SQL. AISQL extends SQL with primitive operators that bring LLM capabilities to the relational model: AI_COMPLETE for text generation, AI_FILTER for semantic filtering, AI_JOIN for semantic joins, AI_CLASSIFY for categorization, and AI_AGG and AI_SUMMARIZE_AGG for semantic aggregations and text summarizations, respectively. These operators naturally compose with traditional SQL constructs, allowing users to write declarative queries that blend relational operations with semantic reasoning. For example, a product manager analyzing customer feedback can filter support transcripts for conversations where customers expressed frustration (AI_FILTER), semantically join them with a product catalog to identify which specific products were discussed (AI_JOIN), classify the severity of each issue (AI_CLASSIFY), and generate executive summaries grouped by product category (AI_SUMMARIZE_AGG) – all within a single declarative SQL query that combines structured sales data with unstructured conversational text.
@@
AISQL builds on an emerging body of work in semantic query processing. Systems such as LOTUS (patel2025semantic) introduced semantic operators for processing pandas-like dataframes, Palimpzest (liu2025palimpzest) provided a declarative framework for LLM-powered data transformations, while ThalamusDB (jo2025thalamusdb) explored multimodal query processing with natural language predicates. Commercial systems have also adopted this direction: Google BigQuery (fernandes2015bigquery) provides AI functions such as AI.CLASSIFY for semantic operations, while Azure SQL exposes row-level LLM invocations within SQL queries. However, making semantic operators efficient at production scale in a distributed database remains a fundamental challenge. Such operators are orders of magnitude more expensive than traditional SQL operations. For example, a single AI_FILTER predicate applied to a million-row table could invoke an LLM thousands of times, leading to prohibitive execution costs and latency. Traditional query optimization heuristics, such as minimizing join costs or pushing filters below joins, can produce extremely inefficient plans when AI operators are involved. Furthermore, semantic operators behave as black boxes to the query optimizer. Unlike conventional predicates, their cost and selectivity cannot be inferred from historical statistics or column distributions. Estimating how many rows will pass an AI-based filter, or how expensive that evaluation will be, is nearly impossible at compile time. Yet, these placement and cost estimates can alter overall query performance by orders of magnitude.
@@
The AISQL query execution engine addresses these challenges through a combination of novel optimization techniques and runtime adaptations, showcasing that semantic operations can be effectively integrated into database engines, thus enabling use cases that analyze both structured and unstructured data. Production deployment data from Snowflake customers directly shaped our design choices. We observed that LLM inference costs far exceed traditional query processing expenses, with AI operations accounting for the dominant share of execution budgets. Multi-table queries are also prevalent, representing nearly 40% of all AISQL workloads and consuming well over half of aggregate computation time (see Section 4 for details). These observations guide the design of our optimization strategies that are applied at both the query and individual operator level:
@@
AI-aware query optimization. The AISQL optimizer treats AI inference cost as a first-class optimization objective. Rather than optimizing solely for traditional metrics like join cardinality, the optimizer considers the monetary and computational cost of LLM invocations when determining operator placement and predicate evaluation order. For instance, the optimizer may pull expensive multi-modal AI_FILTER predicates above joins or reorder multiple AI predicates based on their relative cost, even when traditional selectivity-based heuristics would suggest otherwise. These optimizations enhance the performance of the AISQL query, with the degree of improvement depending on factors such as the selectivities of various query operators (e.g., predicates, joins). Across synthetic datasets and queries modeled after actual Snowflake customer workloads, we have measured 2–8 more performant execution plans. During execution, runtime statistics about predicate cost and selectivity enable dynamic reordering to adapt to actual data distributions. By reasoning about LLM cost directly during query planning, AISQL now makes even complex multi-modal workloads efficient and scalable.
@@
Adaptive model cascades. To reduce inference costs and maintain query prediction accuracy, AISQL implements model cascades for AI_FILTER operations. A fast and inexpensive proxy model processes most rows, while a powerful oracle model handles only uncertain cases. The system learns confidence thresholds online through importance sampling and adaptive threshold estimation, partitioning rows into accept, reject, and uncertainty regions. Our algorithm is tailored for a distributed execution environment and achieves 2–6 speedups across various benchmarks while maintaining 90–95% of oracle model quality. On the NQ dataset, the cascade achieves 5.85 speedup and preserves the near-identical F1 score; see Section 6.2 for details.
@@
Query rewriting for semantic joins. Naive semantic joins using AI_FILTER typically require a quadratic number of inference calls making them impractical for large datasets. The AISQL engine rewrites certain join patterns into multi-label classification problems using AI_CLASSIFY, reducing complexity from quadratic to linear. Our implementation not only achieves 15–70 speedups across a number of datasets but often improves prediction quality by enabling better comparative reasoning. On the CNN dataset, the rewrite achieves 69.52 speedup, reducing execution time from 4.4 hours to 3.8 minutes; see Section 6.3 for details.
@@
AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding. It is fully integrated into the SQL engine and is optimized for efficiency and high prediction quality, as demonstrated by our extensive experimental analysis.
The remainder of this paper is organized as follows. Section 2 describes Snowflake’s architecture with the Cortex Platform for AI workloads. Section 3 introduces AISQL’s operator semantics and SQL interface. Section 4 analyzes customer workload characteristics. Section 5 presents the query execution engine, including AI-aware query optimization, adaptive model cascades, and query rewriting for semantic joins. Section 6 provides an experimental evaluation. Section 7 discusses related work. We conclude in Section 8.
@@ 2. Architecture
In this section, we describe how Snowflake’s architecture has evolved to support AI workloads. Separation of compute and storage is a key design decision in Snowflake’s architecture that lets users scale query-processing resources (i.e., virtual warehouses) independently of the amount of data stored in Snowflake tables. The Cloud Services layer handles authentication, query compilation, and coordination of SQL execution.
@@
The Inference Engines provide the core capabilities of the Cortex Platform. Each engine is a specialized service that runs trained models (e.g., LLMs) to generate predictions or insights from data. An engine manages both the underlying hardware (e.g., GPUs) and the inference stack (e.g., vLLM). The Cortex Platform automatically scales engines up or down to match fluctuations in inference demand.
@@
The Cortex Platform supports a broad set of models from partners including OpenAI, Anthropic, and Meta, enabling users to build and evaluate AI applications easily. For each inference request, the Cortex Platform determines whether to execute it using one of its Inference Engines or forward it to a partner endpoint. For instance, requests for GPT models are routed to OpenAI’s endpoints.
@@ 3. AISQL Operators
In this section, we introduce the core operators that power semantic data processing in Cortex AISQL. These operators extend SQL with natural-language semantics, allowing users to express complex semantic tasks directly within queries. For each operator, we specify its signature, behavior, and typical usage, and show how the operators compose with standard SQL while exposing data-parallel execution to support rich analytical workflows. Table 1 shows the definitions of the key semantic operators implemented in AISQL.
@@ 3.1. Map
AISQL supports simple map or projection operations using the operator. Users can use to transform text expressions using a low-level interface to an LLM and receive a text response for each row provided based on the task instruction . In addition, a PROMPT object provides an ease-of-use interface to include image or other multimodal data to the operator.
@@ 3.2. Filter
is a boolean operator, enabling usage in a SQL WHERE clause. For example, users can filter the rows of a table using a natural language predicate as follows:
@@ 3.4. Classify
This function generalizes beyond sentiment analysis to any classification problem expressible in natural language—such as intent detection, topic tagging, severity rating, or customer status identification—while maintaining compatibility with SQL’s composable semantics.
@@ 3.5. Reduce
As can be seen in the pseudo-code of Algorithm 1, the aggregate function iterates over rows, accumulating text rows in a row buffer . Once the row buffer has exceed the BATCH_SIZE token limit, generates intermediate states containing the most important information for summarization, which are all inserted into an intermediate state buffer .
@@
Finally, after all rows are processed, LLM.Extract and LLM.Combine are executed until and . The final user-facing summary is generated by an LLM invocation to
@@ 3.5.1. Task Specific Aggregation
Algorithmically, is nearly identical to AI_SUMMARIZE_AGG, excepting that each LLM request to LLM.Extract, LLM.Combine, and LLM.Summarize support an additional argument for a user-provided natural language task instruction .
@@ 4. Customer Workloads
As discussed earlier, executing semantic operations inside a database engine introduces several system-design challenges. At the same time, customers are adapting to the AI-centric data management era and explore new workloads and usage patterns to get the most value out of AI. This section analyzes AISQL workloads executed over a three-month period (July–September 2025). We considered actual Snowflake customer workloads across multiple deployments and cloud providers. We report observations intended to inform practitioners and motivate further research. Our main observations from the workload analysis are as follows:
AI operators dominate AISQL query cost. This observation follows from the analysis of AISQL query composition by statement type (see Figure 2) and the cost distribution between LLM GPU credits and SQL Warehouse credits for each statement type (see Figure 4). Thus, in the next section, we present optimization techniques that reduce the cost of running AISQL queries with (a) AI-Aware Query Optimization that takes into account the high cost of executing AI operators to create an effective query plan (see Section 5.1); and (b) Adaptive Model Cascades that uses a lightweight proxy LLM to process most rows while using a larger oracle LLM to handle difficult cases where the proxy model is uncertain (see Section 5.2).
@@ 5. AISQL Query Execution Engine
In this section, we describe the modifications made to Snowflake’s core database engine to enable AISQL queries, i.e. SQL queries that use one or more AI operators. Section 5.1 presents query plan optimizations tailored to AISQL workloads. Section 5.2 discusses the use of model cascade techniques to improve the efficiency of the AI_FILTER operator. Section 5.3 details an optimization that rewrites semantic joins into a multi-label classification problem using AI_CLASSIFY, and Section 5.4 covers enhancements applied to aggregation operators such as AI_AGG and AI_SUMMARIZE_AGG.
@@ 5.1. Optimizing AI Operators
One of the key challenges in optimizing AISQL queries is that AI operators are typically treated as black boxes by the optimizer. For example, without historical data from previous invocations on the same data, it is not possible to estimate the selectivity and hardware cost of an AI_FILTER predicate. At the same time, the per-row monetary and runtime costs of AI operators are typically orders of magnitude higher than those of other typical SQL operators and functions. Consequently, the suboptimal placement of an AI operator in the execution plan can have devastating effects on both cost and execution time. This problem is not new in database systems. Optimizing user-defined functions (UDFs) in database engines shares many similarities with the problem of optimizing AI operators, and many of the techniques published 20 years ago remain relevant (hellerstein6caching; hellestein93predicate; chaudhuri99optimization).
@@
To better illustrate the challenges related to optimizing AISQL queries, let us consider the following scenario: a web application like arxiv.org manages research papers and allows users to search for papers using semantic operations. For simplicity, we assume that the content of each research paper is stored in two relational tables papers and paper_images. The former stores the basic information for each paper, such as ”title” and ”authors”, whereas the latter stores information about extracted images and graphs for each paper. Both tables have a column of type FILE that links each row with a file stored in cloud storage (e.g., AWS S3). A simplified schema and some sample data are shown in Figure 6.
@@
Now let us consider a user who is interested in identifying the research papers that: a) were published in a specific time period, b) discuss ”energy efficiency in database systems”, and c) include a figure with results using the well-known TPC-H benchmark. Finally, the user is interested in compiling a summary of the corresponding abstracts. This query requires inspecting both the textual and image contents of research papers. Today, the user can write the following AISQL query in Snowflake as follows:
@@
The AISQL query in this example has three predicates with different cost and performance characteristics. The first is a predicate on a ”date” column, which is relatively ”cheap” to evaluate and for which most database engines can do a reasonably good job at estimating its selectivity. The second predicate is an AI_FILTER on a VARCHAR column, which is more expensive as it invokes an LLM for each row and its selectivity is unknown at compile time. The execution cost can be estimated based on the average number of tokens in the column values. The third predicate is also an AI_FILTER, but it is applied to images and invokes a multimodal model. Multimodal models are typically larger and more expensive than models operating only on text input.
@@
Suppose the cost of each predicate is ignored. In that case, the optimizer may generate Plan A that is shown in Figure 7, in which all predicates are ”pushed” below the join, with the underlying assumption that join is an expensive operator and hence the number of rows sent to this operator should be minimized. In addition, the most selective predicate (AI_FILTER in this example) is applied first on the rows scanned from the ”papers” table. Plan A results in 110,000 LLM calls; the numbers in the parentheses indicate the output cardinality of each operator.
@@
Figure 7 shows a better execution plan (Plan B) that takes into account the cost of AI predicates and tries to minimize the number of LLM calls. By ”pulling” the image AI_FILTER predicate above the join and by changing the order in which predicates are applied on the rows from the ”papers” table, the compiler generates an execution plan that results in 330 LLM calls, which results in a 300 improvement in both cost and execution time!
@@
This example highlights two of the optimizations that the Snowflake execution engine is applying to improve the performance of AISQL queries (see Section 6 for an experimental evaluation), namely: a) predicate reordering, b) optimizing query plans for AI inference cost. Predicate reordering has both a compiler and an execution component. During compilation, the predicate evaluation order is determined based on the relative cost of each predicate, i.e., the most expensive predicates are evaluated last. During execution, runtime statistics are collected on the cost and selectivity of predicates. These statistics are used to determine at runtime whether a different predicate evaluation order is more effective. For instance, if an AISQL query evaluates two AI_FILTER predicates on text columns from the same input, we can determine at runtime which one is more selective and change the evaluation order so that the more selective AI_FILTER predicate is applied first.
@@
As we discussed in Section 4, AISQL queries typically spend a significant fraction of execution time in AI operators. Consequently, optimizing AISQL queries requires us to rethink the optimization criteria used in the compiler. We initially considered how AI_FILTER predicates are placed in the execution plan with respect to joins and aggregations, as this problem showed up in numerous customer workloads. Although the selectivity of an AI_FILTER is unknown at compile time, we identified that simply optimizing the total number of AI inference calls produces good results in many cases. However, as AISQL queries become more complex and compiler estimations less reliable, these kinds of optimizations become less effective in practice. To that end, we are considering dynamic optimization techniques and caching runtime statistics that can help the compiler make more informed decisions in subsequent AISQL queries. Additionally, we are exploring methods to generalize the compiler’s optimization of AI functions and operator placement in complex execution plans.

### v3

@@ Abstract.
Integrating AI’s semantic capabilities directly into SQL is a key goal, as it would enable users to write declarative queries that blend relational operations with semantic reasoning over both structured and unstructured data. However, achieving production-scale efficiency for semantic operations presents significant hurdles. Semantic operations inherently cost more than traditional SQL operations and possess different latency and throughput characteristics, making their application at large scales prohibitively expensive. Current query execution engines are not designed to optimize semantic operations, often leading to suboptimal query plans. The challenge is compounded because the cost and selectivity of semantic operations are typically unknown at compilation time. Snowflake’s Cortex AISQL is a query execution engine that addresses these challenges through three novel techniques informed by production deployment data from Snowflake customers. First, AI-aware query optimization treats AI inference cost as a first-class optimization objective, reasoning about large language model (LLM) cost directly during query planning to achieve 2–8 speedups. Second, adaptive model cascades reduce inference costs by routing most rows through a fast proxy model while escalating uncertain cases to a powerful oracle model, with 2–6 speedups at 90–95% of oracle model quality. Third, semantic join query rewriting lowers the quadratic time complexity of join operations to linear through reformulation as multi-label classification tasks, for 15–70 speedups, often with improved prediction quality. AISQL is deployed in production at Snowflake, where it powers diverse customer workloads across analytics, search, and content understanding.
@@ 1. Introduction
The growth of large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Wei et al., 2022) has transformed how organizations interact with data. Modern enterprises now store vast volumes of unstructured content such as documents, images, audio, and text alongside traditional structured data in relational tables. While contemporary data warehouses excel at processing structured data with SQL, they lack native support for semantic reasoning over unstructured information. As a result, users who need to filter documents by topic relevance, join tables based on semantic similarity, or summarize customer feedback must export data to external systems, write custom scripts, or build complex pipelines that orchestrate LLM APIs outside the database. The friction is significant: data must be moved across systems, increasing latency and cost; manual orchestration becomes error-prone; and the database loses the ability to optimize the workload end-to-end.
@@
We present Snowflake’s Cortex AISQL, a SQL engine that enables users to talk to their data, both structured and unstructured, through native AI operators integrated directly into SQL. AISQL extends SQL with primitive operators that bring LLM capabilities to the relational model: AI_COMPLETE for text generation, AI_FILTER for semantic filtering, AI_JOIN for semantic joins, AI_CLASSIFY for categorization, and AI_AGG and AI_SUMMARIZE_AGG for semantic aggregations and text summarizations, respectively. These operators naturally compose with traditional SQL constructs, allowing users to write declarative queries that blend relational operations with semantic reasoning. For example, a product manager analyzing customer feedback can filter support transcripts for conversations where customers expressed frustration (AI_FILTER), semantically join them with a product catalog to identify which specific products were discussed (AI_JOIN), classify the severity of each issue (AI_CLASSIFY), and generate executive summaries grouped by product category (AI_SUMMARIZE_AGG), all within a single SQL query that combines structured sales data with unstructured conversational text.
@@
AISQL builds on an emerging body of work in semantic query processing. Systems such as LOTUS (Patel et al., 2025) introduced semantic operators for processing pandas-like dataframes, Palimpzest (Liu et al., 2025) provided a declarative framework for LLM-powered data transformations, while ThalamusDB (Jo and Trummer, 2024) explored multimodal query processing with natural language predicates. Commercial systems have also adopted this direction: Google BigQuery (Fernandes and Bernardino, 2015) provides AI functions such as AI.CLASSIFY for semantic operations, while Azure SQL exposes row-level LLM invocations within SQL queries (Microsoft, 2025). However, making semantic operators efficient at production scale in a distributed database remains a fundamental challenge. Such operators are orders of magnitude more expensive than traditional SQL operations. For example, a single AI_FILTER predicate applied to a million-row table could invoke an LLM once per row, leading to prohibitive execution costs and latency. Traditional query optimization heuristics, such as minimizing join costs or pushing filters below joins, can produce extremely inefficient plans when AI operators are involved. Semantic operators also behave as black boxes to the query optimizer. Unlike conventional predicates, their cost and selectivity cannot be inferred from historical statistics or column distributions. Estimating how many rows will pass an AI-based filter, or how expensive that evaluation will be, is nearly impossible at compile time. Yet, these placement and cost estimates can alter overall query performance dramatically.
@@
The AISQL query execution engine addresses these challenges through novel optimization techniques and runtime adaptations shaped by production deployment data from Snowflake customers. LLM inference costs far exceed traditional query processing expenses, with AI operations accounting for the dominant share of execution budgets. Multi-table queries are also prevalent, representing nearly 40% of all AISQL workloads and consuming over half of aggregate computation time (see Section 4). These observations guide our optimization strategies applied at both the query and individual operator level:
@@
AI-aware query optimization. The AISQL optimizer treats AI inference cost as a first-class optimization objective. Rather than optimizing solely for traditional metrics like join cardinality, the optimizer considers the monetary and computational cost of LLM invocations when determining operator placement and predicate evaluation order. For instance, the optimizer may pull expensive multi-modal AI_FILTER predicates above joins or reorder multiple AI predicates based on their relative cost, even when traditional selectivity-based heuristics would suggest otherwise. Across synthetic datasets and queries modeled after actual Snowflake customer workloads, we have measured 2–8 faster execution plans. During execution, runtime statistics about predicate cost and selectivity enable dynamic reordering to adapt to actual data distributions.
@@
Adaptive model cascades. To reduce inference costs and maintain query prediction accuracy, AISQL implements model cascades for AI_FILTER operations. A fast and inexpensive proxy model processes most rows, while a powerful oracle model handles only uncertain cases. The system learns confidence thresholds online through importance sampling and adaptive threshold estimation, partitioning rows into accept, reject, and uncertainty regions (Liskowski and Schmaus, 2026). Our algorithm is tailored for a distributed execution environment and achieves 2–6 speedups across various benchmarks while maintaining 90–95% of oracle model quality. On the NQ dataset, the cascade achieves 5.85 speedup and preserves a near-identical F1 score; see Section 6.2 for details.
@@
Query rewriting for semantic joins. Naive semantic joins using AI_FILTER typically require a quadratic number of inference calls, making them impractical for large datasets. The AISQL engine rewrites certain join patterns into multi-label classification problems using AI_CLASSIFY and reduces complexity from quadratic to linear. Our implementation not only achieves 15–70 speedups across a number of datasets but often improves prediction quality because the multi-label formulation enables better comparative reasoning. On the CNN dataset, the rewrite achieves 69.52 speedup, reducing execution time from 4.4 hours to 3.8 minutes; see Section 6.3 for details.
@@
AISQL is deployed in production at Snowflake, fully integrated into the SQL engine, where it powers diverse customer workloads across analytics, search, and content understanding.
@@ 2. Architecture
Snowflake’s architecture has evolved to support AI workloads. Separation of compute and storage is a key design decision in Snowflake’s architecture that lets users scale query-processing resources (i.e., virtual warehouses) independently of the amount of data stored in Snowflake tables. The Cloud Services layer handles authentication, query compilation, and coordination of SQL execution.
@@
Each Inference Engine is a specialized service that hosts open-weight models (e.g., Llama, Mistral) on Snowflake-managed GPU infrastructure to generate predictions or insights from data. An engine manages both the underlying hardware and the inference stack (e.g., vLLM), distinct from the partner endpoints described below. The Cortex Platform automatically scales engines up or down to match fluctuations in inference demand.
@@
The Cortex Platform supports a broad set of models from partners including OpenAI, Anthropic, and Meta. For each inference request, the Cortex Platform determines whether to execute it using one of its Inference Engines or forward it to a partner endpoint. Requests for GPT models, for example, are routed to OpenAI’s endpoints.
@@ 3. AISQL Operators
The core operators that power semantic data processing in Cortex AISQL extend SQL with natural-language semantics, allowing users to express complex analytical tasks directly within queries. For each operator, we specify its signature, behavior, and typical usage, and show how the operators compose with standard SQL while exposing data-parallel execution to support rich analytical workflows. Table 1 shows the definitions of the key semantic operators implemented in AISQL.
@@ 3.1. Map
AISQL supports simple map or projection operations using the operator. Users can use to transform text expressions using a low-level interface to an LLM and receive a text response for each row provided based on the task instruction . A PROMPT object provides a convenience interface for including image or other multimodal data in the operator call.
@@ 3.2. Filter
is a boolean operator for use in a SQL WHERE clause. For example, users can filter the rows of a table using a natural language predicate as follows:
@@ 3.4. Classify
The operator generalizes beyond sentiment analysis to any classification problem expressible in natural language (intent detection, topic tagging, severity rating, or customer status identification) while maintaining compatibility with SQL’s composable semantics.
@@ 3.5. Reduce
The pseudo-code in Algorithm 1 shows how the aggregate function iterates over rows, accumulating text rows in a row buffer . Once the row buffer has exceed the BATCH_SIZE token limit, generates intermediate states containing the most important information for summarization, which are all inserted into an intermediate state buffer .
@@
Finally, after all rows are processed, LLM.Extract and LLM.Combine are executed until and . The final user-facing summary is generated by an LLM invocation to .
@@ 3.5.1. Task Specific Aggregation
Algorithmically, is nearly identical to AI_SUMMARIZE_AGG, except that each LLM request to LLM.Extract, LLM.Combine, and LLM.Summarize supports an additional argument for a user-provided natural language task instruction .
@@ 4. Customer Workloads
As discussed earlier, executing semantic operations inside a database engine introduces several system-design challenges. At the same time, customers are adapting to the AI-centric data management era and exploring new workloads and usage patterns to extract maximum value from AI. We analyze AISQL workloads executed over a three-month period (July–September 2025) across multiple Snowflake deployments and cloud providers. The observations below are intended to inform practitioners and motivate further research. Our main observations from the workload analysis are as follows:
AI operators dominate AISQL query cost. The observation follows from the analysis of AISQL query composition by statement type (see Figure 2) and the cost distribution between LLM GPU credits and SQL Warehouse credits for each statement type (see Figure 4). Thus, in the next section, we present optimization techniques that reduce the cost of running AISQL queries with (a) AI-Aware Query Optimization that takes into account the high cost of executing AI operators to create an effective query plan (see Section 5.1); and (b) Adaptive Model Cascades that uses a lightweight proxy LLM to process most rows while using a larger oracle LLM to handle difficult cases where the proxy model is uncertain (see Section 5.2).
@@ 5. AISQL Query Execution Engine
Snowflake’s core database engine required several modifications to support AISQL queries, i.e. SQL queries that use one or more AI operators. Section 5.1 presents query plan optimizations tailored to AISQL workloads. Section 5.2 discusses the use of model cascade techniques to improve the efficiency of the AI_FILTER operator. Section 5.3 details an optimization that rewrites semantic joins into a multi-label classification problem using AI_CLASSIFY, and Section 5.4 covers enhancements applied to aggregation operators such as AI_AGG and AI_SUMMARIZE_AGG. Predicate reordering and cost-based placement adapt well-known optimization techniques for expensive predicates (Hellerstein and Naughton, 1996; Hellerstein and Stonebraker, 1993; Chaudhuri and Shim, 1999) to the AISQL setting. The remaining two techniques are novel contributions: adaptive model cascades introduce a streaming threshold-learning algorithm with formal precision-recall guarantees (Liskowski and Schmaus, 2026), and the join-to-classification rewrite exploits the structure of semantic join predicates to reduce quadratic-cost joins to linear classification.
@@ 5.1. Optimizing AI Operators
One of the key challenges in optimizing AISQL queries is that AI operators are treated as black boxes by the optimizer. For example, without historical data from previous invocations on the same data, it is not possible to estimate the selectivity and hardware cost of an AI_FILTER predicate. At the same time, the per-row monetary and runtime costs of AI operators are orders of magnitude higher than those of conventional SQL operators and functions. Consequently, the suboptimal placement of an AI operator in the execution plan can have a dramatic impact on both cost and execution time. The problem is not new in database systems. Optimizing user-defined functions (UDFs) in database engines shares many similarities with the problem of optimizing AI operators, and many of the techniques published 20 years ago remain relevant (Hellerstein and Naughton, 1996; Hellerstein and Stonebraker, 1993; Chaudhuri and Shim, 1999).
@@
To illustrate the challenges of optimizing AISQL queries, let us consider the following scenario: a web application like arxiv.org manages research papers and allows users to search for papers using semantic operations. For simplicity, we assume that the content of each research paper is stored in two relational tables papers and paper_images. The former stores the basic information for each paper, such as ”title” and ”authors”, whereas the latter stores information about extracted images and graphs for each paper. Both tables have a column of type FILE that links each row with a file stored in cloud storage (e.g., AWS S3). A simplified schema and some sample data are shown in Figure 6.
@@
Now let us consider a user who is interested in identifying the research papers that: a) were published in a specific time period, b) discuss ”energy efficiency in database systems”, and c) include a figure with results using the TPC-H benchmark. Finally, the user is interested in compiling a summary of the corresponding abstracts. This query requires inspecting both the textual and image contents of research papers. Today, the user can write the following AISQL query in Snowflake:
@@
The AISQL query in this example has three predicates with different cost and performance characteristics. The first is a predicate on a ”date” column, which is relatively ”cheap” to evaluate and for which most database engines can reasonably estimate its selectivity. The second predicate is an AI_FILTER on a VARCHAR column, which is more expensive as it invokes an LLM for each row and its selectivity is unknown at compile time. The execution cost can be estimated based on the average number of tokens in the column values. The third predicate is also an AI_FILTER, but it is applied to images and invokes a multimodal model. Multimodal models are generally larger and more expensive than models operating only on text input.
@@
Suppose the cost of each predicate is ignored. In that case, the optimizer may generate Plan A, shown in Figure 7, in which all predicates are ”pushed” below the join, with the underlying assumption that join is an expensive operator and hence the number of rows sent to this operator should be minimized. The most selective predicate (AI_FILTER in this example) is applied first on the rows scanned from the ”papers” table. Plan A results in 110,000 LLM calls; the numbers in the parentheses indicate the output cardinality of each operator.
@@
Figure 7 shows a better execution plan (Plan B) that takes into account the cost of AI predicates and tries to minimize the number of LLM calls. By ”pulling” the image AI_FILTER predicate above the join and by changing the order in which predicates are applied on the rows from the ”papers” table, the compiler generates an execution plan with only 330 LLM calls, a 300 improvement in both cost and execution time.
@@
The example highlights two of the optimizations that the Snowflake execution engine applies to improve the performance of AISQL queries (see Section 6 for an experimental evaluation), namely: a) predicate reordering, b) optimizing query plans for AI inference cost. Predicate reordering has both a compiler and an execution component. During compilation, the predicate evaluation order is determined based on the relative cost of each predicate, i.e., the most expensive predicates are evaluated last. During execution, runtime statistics are collected on the cost and selectivity of predicates. These statistics are used to determine at runtime whether a different predicate evaluation order is more effective. For instance, if an AISQL query evaluates two AI_FILTER predicates on text columns from the same input, we can determine at runtime which one is more selective and change the evaluation order so that the more selective AI_FILTER predicate is applied first.
@@
As we discussed in Section 4, AISQL queries spend a significant fraction of execution time in AI operators. Consequently, optimizing AISQL queries requires us to rethink the optimization criteria used in the compiler. We initially considered how AI_FILTER predicates are placed in the execution plan with respect to joins and aggregations, as this problem showed up in numerous customer workloads. Although the selectivity of an AI_FILTER is unknown at compile time, we identified that simply optimizing the total number of AI inference calls produces good results in many cases. However, as AISQL queries become more complex and compiler estimations less reliable, such optimizations become less effective in practice. To that end, we are considering dynamic optimization techniques, caching runtime statistics, and methods to generalize AI function placement in complex execution plans.